# How to Train Your LLM Web Agent: A Statistical Diagnosis

### v3

@@ Appendix I Random Search Space
The most important attribute for the robustness of the boostrap is its ability to exploit the Central Limit Theorem (CLT) to approximate the sampling distribution of a statistic. Using the CLT it can allow us to estimate its expectation and variability without parametric assumptions [12].
@@
This follows from the law of large numbers [24], assuming sufficient coverage of other hyperparameters $\mathcal{G}$. The rate of convergence depends on the tail behavior of the conditional distribution $M\mid H=h$. Since our evaluation metric is bounded in $[0,1]$, the sub-Gaussian assumption holds automatically by Hoeffding’s Lemma [14], ensuring concentration of the estimate. Under these conditions, the expected bias decays as $\mathcal{O}(1/n_{h})$, where $n_{h}$ is the number of runs with $H=h$.
These are well-known results, which is why the bootstrap is an extremely useful and simple algorithm for estimating any statistic of interest [11, 12].
@@
We evaluate generalization by training only on the train split and reporting performance on held-out tasks from the test split. For the held-out goals metric, we instantiate goal variations using seed ranges $[0,2]$ [8] for training tasks (3 goals per task) and $[0,9]$ [8] for test tasks (10 goals per task). Below we list the exact task identifiers used in our experiments for both benchmarks. Task names are the registry keys from the respective environments.

### v4

@@ Appendix I Random Search Space
The most important attribute for the robustness of the boostrap is its ability to exploit the Central Limit Theorem (CLT) to approximate the sampling distribution of a statistic. Using the CLT it can allow us to estimate its expectation and variability without parametric assumptions (Efron1994).
@@
This follows from the law of large numbers (Wainwright2019), assuming sufficient coverage of other hyperparameters $\mathcal{G}$. The rate of convergence depends on the tail behavior of the conditional distribution $M\mid H=h$. Since our evaluation metric is bounded in $[0,1]$, the sub-Gaussian assumption holds automatically by Hoeffding’s Lemma (Hoeffding1963), ensuring concentration of the estimate. Under these conditions, the expected bias decays as $\mathcal{O}(1/n_{h})$, where $n_{h}$ is the number of runs with $H=h$.
These are well-known results, which is why the bootstrap is an extremely useful and simple algorithm for estimating any statistic of interest (Efron1979; Efron1994).
@@
We evaluate generalization by training only on the train split and reporting performance on held-out tasks from the test split. For the held-out goals metric, we instantiate goal variations using seed ranges $[0,2]$ chezelles2025the for training tasks (3 goals per task) and $[0,9]$ chezelles2025the for test tasks (10 goals per task). Below we list the exact task identifiers used in our experiments for both benchmarks. Task names are the registry keys from the respective environments.